Kathi and Robert Hawn had a pottery stand at the annual skipcack craft fair. They sold some of their pottery at the original price of $9.50 for each piece, but later decreased the price by $2. If they sold all 90 pieces and took in $713, find how many pieces they sold at the original price and how many pieces they sold at the reduced price?
Answer by mananth(16946)

let them sell x pieces at $9.50
and y pieces at $7.50
x+y=90
9.5x+7.5y=713
$9.50 x numbers
$7.50 y numbers

1 x + 1 y = 90 .............1
9.5 x + 7.5 y = 713 .............2
Eliminate y
multiply (1)by -7.5
Multiply (2) by 1
-7.5 x -7.5 y = -675
9.5 x 7.5 y = 713
Add the two equations
2 x = 38
/ 2
x = 19
plug value of x in (1)
1 x + 1 y = 90
19 + 1 y = 90
1 y = 90 -19
1 y = 71
y = 71
$9.50 19 numbers
$7.50 71 numbers
m.ananth@hotmail.ca